# Bridging Model Reference Adaptive Control and Data Informativity

### v1

@@ 5 MRAC with Online Informative Data
We aim to design an adaptive law to guarantee the convergence of $\hat{K}$ and $\hat{L}$ in (7) to some (possibly non-unique) solution of the matching equations (6). We first provide necessary and sufficient conditions for existence of a time at which the system data become informative in the sense of Lemma 3 (cf. Section V-A). We then provide necessary and sufficient conditions for convergence to a solution of the matching equations by only relying on data informativity for model reference control (cf. Section V-B). Finally, stability and convergence for the closed-loop signals are studied (cf. Section V-C) and discussions on existing MRAC solutions are given (cf. Section V-D).
@@ 5.1 Necessary and sufficient conditions for existence of informative time
$T^{*}:=\min\{t\mid$ $(U_{-}(t),X(t))\in\mathfrak{B}_{t}\text{ are informative}$ (19)
[…]
@@
$\Phi_{X_{-}}\!(t)$ ={$&X_{-}(t)&&\text{if }0<t\leq T^{*}+1\\ &\!\!\begin{bmatrix}X_{-}(T^{*})&x(t-1)\\ \end{bmatrix}\!\!&&\text{if }t\!>\!T^{*}\!+\!1\,\text{and}\,|x(t)|\leq\sigma\\ &\Phi_{X_{-}}(t-1)&&\text{otherwise},\end{aligned}\right.$
$\Phi_{X_{+}}\!(t)$ ={$&X_{+}(t)&&\text{if }0<t\leq T^{*}+1\\ &\!\!\begin{bmatrix}X_{+}(T^{*})&x(t)\\ \end{bmatrix}\!\qquad\!\!\!\!\ &&\text{if }t\!>\!T^{*}\!+\!1\,\text{and}\,|x(t)|\leq\sigma\\ &\Phi_{X_{+}}(t-1)&&\text{otherwise},\end{aligned}\right.$
@@
where $\sigma\in\mathbb{R}$ is a given upper bound on the norm of the collected state data. The use of data matrices with both past and current data is inspired by the concurrent learning approach [21], where past and current data are concurrently used to update the control gains. Note that to construct the data matrices in (20) and (21), we need to verify at every time $t$ whether $t\leq T^{*}+1$. We emphasize that, even though the exact value of $T^{*}$ may be unknown at time $t$, one can verify whether $t\leq T^{*}+1$ using (11).
@@
$\Theta(t+1)=\left\{\begin{aligned} &\begin{bmatrix}\Theta(t)\\ 0\end{bmatrix}-\gamma\Phi_{X}(t)^{\top}\Delta(t)&&\text{if }0<t\leq T^{*}\\ &\ \Theta(t)-\gamma\Phi_{X}(t)^{\top}\Delta(t)&&\text{if }t>T^{*},\end{aligned}\right.$ (23)
where $20<\gamma<$ 2 and
@@
$&\Phi_{X}(t)\Theta(t)-\begin{bmatrix}I&0\\ A_{\rm m}&B_{\rm m}\end{bmatrix}&&\text{if }0<t\leq T^{*}\\ &\frac{1}{\|\Phi_{X}(t)\|_{F}^{2}}\left(\Phi_{X}(t)\Theta(t)-\begin{bmatrix}I&0\\ A_{\rm m}&B_{\rm m}\end{bmatrix}\right)&&\text{if }t>T^{*}.\end{aligned}\right.$
@@
We first show that if i), ii) and iii) hold, then there exists a $u_{\mathrm{r}}(t)\in\mathbb{R}^{m}$ satisfying (26). To do so, note that iii) implies that $t\leq T^{*}+1$ so $\Phi_{U_{-}}(t)=U_{-}(t)$ and $\Phi_{X_{-}}(t)=X_{-}(t)$. Moreover, ii) implies that $x(t)\in\operatorname{im}X_{-}(t)$. Therefore, it follows from [37, Theorem 1] that there exist a $\xi\in\mathbb{R}^{n}$ and a nonzero $\eta\in\mathbb{R}^{m}$ such that
@@
Lemma 5.7 demonstrates that, with the specific choice of the input, either $T^{*}\leq n+m$ or $T^{*}=\infty$. In the latter case, there are no gains $(K,L)$ such that (6) holds. Therefore, Lemma 5.7 provides an effective method to check whether the matching equations have a solution. Indeed, $T^{*}=\infty$ if (11) does not hold for $t=n+m$. In other words, $T^{*}$ can always be determined after collecting at most $n+m$ data samples.
@@
(Design of the input). Following [37], the input $u_{r}(t)$ in Lemma 5.7 can be designed as follows: first, find $\xi\in\mathbb{R}^{n}$, $\eta\in\mathbb{R}^{m}\setminus\{0\}$ such that (27) holds; then, design $u_{\mathrm{r}}(t)$ such that (28) holds.
@@
(Design of the parameters). In (20) and (21), $\sigma$ can be designed, for instance, as an upper bound of $x_{\rm m}$, which can be determined since the reference model $A_{\rm m},B_{\rm m}$ and the reference input $r\in\ell_{\infty}^{p}$ are given. In (23), the gain has been taken as a scalar $\gamma\in(0,2)$, for simplicity. However, the convergence in (10) can be generalized to a matrix gain $\Gamma=\Gamma^{\top}$ satisfying $0<\Gamma<2I$. As in gradient-based adaptation, a matrix gain allows different adaptation rates along different directions [16].
@@ 5.2 Necessary and sufficient conditions for convergence to a matching solution
$\Phi_{X}(t)\Theta^{*}=\begin{bmatrix}I&0\\ A_{\rm m}&B_{\rm m}\end{bmatrix}$
@@
$V(\tilde{\Theta})=\frac{1}{2\gamma}\operatorname{tr}\left(\tilde{\Theta}^{\top}\tilde{\Theta}\right).$ (34)
@@
$=$ $\frac{1}{2\gamma}\operatorname{tr}\left(\tilde{\Theta}(t+1)^{\top}\tilde{\Theta}(t+1)\right)\!-\!\frac{1}{2\gamma}\operatorname{tr}\left(\tilde{\Theta}(t)^{\top}\tilde{\Theta}(t)\right)$
$=$ $\!-\!\operatorname{tr}\left(\tilde{\Theta}(t)^{\top}\Phi_{X}(t)^{\top}\Delta(t)\right)\!+\!\frac{\gamma}{2}\operatorname{tr}\left(\Delta(t)^{\top}\Phi_{X}(t)\Phi_{X}(t)^{\top}\Delta(t)\right)$
$=$ $\!-\!\|\Phi_{X}(t)\|_{F}^{2}\cdot\operatorname{tr}\left(\Delta(t)^{\top}\Delta(t)\right)$
$\!+\!\frac{\gamma}{2}\operatorname{tr}\left(\Delta(t)^{\top}\Phi_{X}(t)\Phi_{X}(t)^{\top}\Delta(t)\right)$
$=$ $\!-\!\|\Phi_{X}(t)\|_{F}^{2}\cdot\operatorname{tr}\left(\Delta(t)^{\top}\!\left(I\!-\!\frac{\gamma\Phi_{X}(t)\Phi_{X}(t)^{\top}}{2\|\Phi_{X}(t)\|_{F}^{2}}\right)\Delta(t)\right)\!.$
Since $20<\gamma<$ 2, there exists $\varepsilon>0$ such that $1-\frac{\gamma}{2}\geq\epsilon$. We claim that
$I-\frac{\gamma\Phi_{X}(t)\Phi_{X}(t)^{\top}}{2\|\Phi_{X}(t)\|_{F}^{2}}\geq\varepsilon I.$ (35)
@@
$v^{\top}\!\left(I\!-\!\frac{\gamma\Phi_{X}(t)\Phi_{X}(t)^{\top}}{2\|\Phi_{X}(t)\|_{F}^{2}}\right)v$ $\geq|v|^{2}-\frac{\gamma\|\Phi_{X}(t)\|_{2}^{2}}{2\|\Phi_{X}(t)\|_{F}^{2}}|v|^{2}$
$\geq|v|^{2}-\frac{\gamma}{2}|v|^{2}\geq\varepsilon|v|^{2}.$
@@
From (21), there exists a $c\!>\!0$ such that $\operatorname{tr}\!\left(\Phi_{X}(t)^{\top}\Phi_{X}(t)\right)\!\leq\!c$ for all $t>T^{*}$. Therefore,
@@
$\lim\limits_{t\to\infty}\begin{bmatrix}A_{\rm s}+B_{\rm s}\hat{K}(t)&B_{\rm s}\hat{L}(t)\end{bmatrix}$
@@
Remark 5.13.
(A non-conservative condition on data). According to Theorem 5.11, the convergence to a solution of the matching equations is achieved if and only if $T^{*}$ is finite. This, in turn, is equivalent to a specific condition on the collected data at the time instant $T^{*}$, namely,
$\operatorname{im}\begin{bmatrix}I&0\\ A_{\rm m}&B_{\rm m}\end{bmatrix}\subseteq\operatorname{im}\begin{bmatrix}X_{-}(T^{*})\\ X_{+}(T^{*})\end{bmatrix}.$ (40)
In other words, Theorem 5.11 establishes a non-conservative condition on the collected data to achieve convergence of the adaptive gains to a solution of the matching equations.
Theorem 5.11 provides a novel MRAC approach where neither unique system identification nor structural assumptions on the system apart from controllability are required. We refer the reader to Section V-D for a discussion on structural assumptions in existing MRAC solutions, such as square assumptions on $L$ ($p=m$), or $B_{{\rm s}}$ being known, or $B_{{\rm s}}$ being full column rank. Before this, we discuss the boundedness of the closed-loop signals and the convergence of the error.
@@ 5.3 Stability and boundedness of the closed-loop signals
([41]). Consider a linear time-varying autonomous system
@@
([42]). Consider a nonlinear system
@@
Since $\lim_{t\to\infty}A_{\rm s}+B_{\rm s}\hat{K}(t)=A_{\rm m}$ and $A_{\rm m}$ is Schur, [41, Corollary 3] implies that (44) is uniformly exponentially stable. From Definition 5.14, we have that for any $t_{0}\in\mathbb{Z}_{+}$ and $t>t_{0}$, there exist scalars $\rho^{\prime}>0$ and $\lambda^{\prime}\in(0,1)$ such that
@@
Finally, we demonstrate the convergence of the error. According to [42, Sect. 3.2], for the input-to-state stable system (46), there exists a $\mathcal{K}$-function $f^{\prime}$ such that for any $e(0)\in\mathbb{R}^{n}$,
@@ 5.4 Discussion on excitation and structural assumptions in state-of-the-art MRAC
We now discuss the requirements on collected data imposed by the proposed MRAC method as compared to existing state-of-the-art MRAC methods. Without going into full details, we mention that state-of-the-art approaches like concurrent learning or composite MRAC [20, 21, 22, 23, 24, 25] are based on a Lyapunov design in continuous time, based on the error dynamics
@@
with $A_{\rm m}$ the Hurwitz state matrix of the reference model, $B_{\rm s}$ the control matrix of the system, $\psi$ measured signals, and $\tilde{\Lambda}$ a parameter estimation error. In this paper, we have proposed a discrete-time MRAC method. Although a direct comparison to the state-of-the-art is thus not possible, some considerations can still be made based on the analogy between the error dynamics in (46) and (48). To this end, note that all state-of-the-art solutions with the exception of [25] require the input matrix $B_{\rm s}$ to be known, which is not required in the MRAC method in the current paper. The estimation of $B_{\rm s}$ in [25] leveraged the concept of initial excitation to relax the classical persistence of excitation conditions. When written in discrete time, initial excitation requires the existence of $T\in\mathbb{N}$ and $\delta>0$ such that
@@
satisfy the full rank condition (4). Thus, from Lemma 2, we conclude that the requirements imposed on collected data by initial excitation are such that $(A_{\rm s},B_{\rm s})$ can be uniquely identified. Such an observation further illustrates the significance of Proposition 4.2 and Theorem 4.4. Similar to [25], let us assume that the matching equations (6) have a solution and $p=m$. From Proposition 4.2, we know that (49) implies data informativity for model reference control. However, when $B_{\rm m}$ does not have full column rank, Theorem 4.4 implies that the initial excitation (49) is sufficient but not necessary to solve the model reference control problem. This means that initial excitation imposes stronger conditions on collected data, not required in the proposed MRAC approach.
Let us finally note that most approaches [27, 20, 21, 22, 23, 25, 24, 26] in the multi-variable MRAC literature require $B_{\rm s}$ to have full column rank, which in turns guarantees a unique solution $(K,L)$ of the matching equations (6). This is not required in the proposed MRAC method, which remains valid even when an infinite number of solutions of the matching equations (6) exist. This will be illustrated in the numerical example in the next section.

### v2

@@ 5 MRAC with Online Informative Data
We aim to design an adaptive law to guarantee asymptotic convergence of $\hat{K}$ and $\hat{L}$ in (7) to some (possibly non-unique) solution of the matching equations (6). We first provide necessary and sufficient conditions for existence of a time at which the system data become informative in the sense of Lemma 3 (cf. Section V-A). We then provide necessary and sufficient conditions for asymptotic convergence to a solution of the matching equations by only relying on data informativity for model reference control (cf. Section V-B). Finally, stability and convergence for the closed-loop signals are studied (cf. Section V-C) and discussions on existing MRAC solutions are given (cf. Section V-D).
@@ 5.1 Necessary and sufficient conditions for existence of informative time
$T^{*}=\min\{t\mid$ $(U_{-}(t),X(t))\in\mathfrak{B}_{t}\text{ are informative}$ (19)
$\text{ for model reference control}\}.$
@@
$\Phi_{X_{-}}\!(t)$ $\!=\!\left\{\begin{aligned} &X_{-}(t)&&\text{if }0<t\leq T^{*}+1\\ &\!\!\begin{bmatrix}X_{-}(T^{*})&x(t-1)\\ \end{bmatrix}\!\!&&\text{if }t\!>\!T^{*}\!+\!1\,\text{and}\,|x(t)|\leq\sigma\\ &\Phi_{X_{-}}(t-1)&&\text{otherwise},\end{aligned}\right$.
$\Phi_{X_{+}}\!(t)$ $\!=\!\left\{\begin{aligned} &X_{+}(t)&&\text{if }0<t\leq T^{*}+1\\ &\!\!\begin{bmatrix}X_{+}(T^{*})&x(t)\\ \end{bmatrix}\!\qquad\!\!\!\!\ &&\text{if }t\!>\!T^{*}\!+\!1\,\text{and}\,|x(t)|\leq\sigma\\ &\Phi_{X_{+}}(t-1)&&\text{otherwise},\end{aligned}\right$.
@@
where $\sigma\!\in\!\mathbb{R}$ is a given upper bound on the norm of the collected state data. The use of data matrices with past and current data is inspired by the concurrent learning approach [chowdhary2010concurrent], where past and current data are concurrently used to update the control gains. Note that to construct the data matrices in (20) and (21), we need to verify at every time $t$ whether $t\leq T^{*}+1$. We emphasize that, even though the exact value of $T^{*}$ may be unknown at time $t$, one can verify whether $t\leq T^{*}+1$ via the linear algebraic condition (11).
@@
$\Theta(t+1)=\left\{\begin{aligned} &\begin{bmatrix}\Theta(t)-\gamma\Phi_{X}(t)^{\top}\Delta(t)\\ 0\end{bmatrix}&&\text{if }0<t\leq T^{*}\\ &\ \Theta(t)-\gamma\Phi_{X}(t)^{\top}\Delta(t)&&\text{if }t>T^{*},\end{aligned}\right$. (23)
where $0<\gamma<2$ and
@@
$\left\{\begin{aligned} &\Phi_{X}(t)\Theta(t)-\begin{bmatrix}I&0\\ A_{\rm m}&B_{\rm m}\end{bmatrix}&&\text{if }0<t\leq T^{*}\\ &\frac{1}{\|\Phi_{X}(t)\|_{F}^{2}}\left(\Phi_{X}(t)\Theta(t)-\begin{bmatrix}I&0\\ A_{\rm m}&B_{\rm m}\end{bmatrix}\right)&&\text{if }t>T^{*}.\end{aligned}\right$.
@@
We first show that if i), ii) and iii) hold, then there exists a $u_{\mathrm{r}}(t)\in\mathbb{R}^{m}$ satisfying (26). To do so, note that iii) implies that $t\leq T^{*}+1$ so $\Phi_{U_{-}}(t)=U_{-}(t)$ and $\Phi_{X_{-}}(t)=X_{-}(t)$. Moreover, ii) implies that $x(t)\in\operatorname{im}X_{-}(t)$. Therefore, it follows from [van2021beyond, Theorem 1] that there exist a $\xi\in\mathbb{R}^{n}$ and a nonzero $\eta\in\mathbb{R}^{m}$ such that
@@
Lemma 5.7 demonstrates that, with the specific choice of the input, either $T^{*}\leq n+m$ or $T^{*}=\infty$. In the latter case, there are no gains $(K,L)$ such that (6) holds. Therefore, Lemma 5.7 provides an effective method to check whether the matching equations (6) have a solution. Indeed, $T^{*}=\infty$ if (11) does not hold for $t=n+m$. In other words, $T^{*}$ can always be determined after collecting at most $n+m$ data samples.
@@
(Determining $T^{*}$). To determine $T^{*}$, one needs to verify the linear algebraic condition (11). Note that (11) is equivalent to
[…]
which provides a computationally efficient way to verify (11). Furthermore, (32) implies that $n+\operatorname{rank}(B_{\rm m})\leq T^{*}\leq n+m$, so it suffices to check (32) only for $t\geq n+\operatorname{rank}(B_{\rm m})$.
@@
(Design of the input). Following [van2021beyond], the input $u_{r}(t)$ in Lemma 5.7 can be designed as follows: first, find $\xi\in\mathbb{R}^{n}$, $\eta\in\mathbb{R}^{m}\setminus\{0\}$ such that (27) holds; then, design $u_{\mathrm{r}}(t)$ such that (28) holds.
Remark 5.11.
(Design of the parameters). In (20) and (21), $\sigma$ can be designed, for instance, as an upper bound of $x_{\rm m}$, which can be determined since the reference model $A_{\rm m},B_{\rm m}$ and the reference input $r\in\ell_{\infty}^{p}$ are given. In (23), the gain has been taken as a scalar $\gamma\in(0,2)$, for simplicity [^2] Requirements analogous to $\gamma\in(0,2)$ are common in discrete-time gradient-based adaptation [ioannou2006adaptive, Lemma 4.1.1]. It will be shown in the proof of Theorem 5.12 that this choice of $\gamma$ also ensures the existence of a limit of the Lyapunov function, which in turn guarantees (10). . However, the scalar $\gamma$ can be generalized to a matrix gain $\Gamma=\Gamma^{\top}$ satisfying $0<\Gamma<2I$, which allows different adaptation rates along different directions [ioannou2006adaptive].
@@ 5.2 Necessary and sufficient conditions for convergence to a matching solution
$\Phi_{X}(t)\Theta^{*}=\begin{bmatrix}I&0\\ A_{\rm m}&B_{\rm m}\end{bmatrix}$ (34)
@@
[…]
@@
$=$ $\frac{1}{2}\operatorname{tr}\left(\tilde{\Theta}(t+1)^{\top}\tilde{\Theta}(t+1)\right)-\frac{1}{2}\operatorname{tr}\left(\tilde{\Theta}(t)^{\top}\tilde{\Theta}(t)\right)$
$=$ $-\gamma\operatorname{tr}\left(\tilde{\Theta}(t)^{\top}\Phi_{X}(t)^{\top}\Delta(t)\right)$
$+\frac{\gamma^{2}}{2}\operatorname{tr}\left(\Delta(t)^{\top}\Phi_{X}(t)\Phi_{X}(t)^{\top}\Delta(t)\right)$
$=$ $-\gamma\|\Phi_{X}(t)\|_{F}^{2}\cdot\operatorname{tr}\left(\Delta(t)^{\top}\Delta(t)\right)$
$+\frac{\gamma^{2}}{2}\operatorname{tr}\left(\Delta(t)^{\top}\Phi_{X}(t)\Phi_{X}(t)^{\top}\Delta(t)\right)$
$=$ $-\|\Phi_{X}(t)\|_{F}^{2}\cdot\operatorname{tr}\left(\Delta(t)^{\top}\left(\gamma I-\frac{\gamma^{2}\Phi_{X}(t)\Phi_{X}(t)^{\top}}{2\|\Phi_{X}(t)\|_{F}^{2}}\right)\Delta(t)\right)\!.$
Since $0<\gamma<2$, there exists $\varepsilon>0$ such that $\gamma-\frac{\gamma^{2}}{2}\geq\varepsilon$. We claim that
$\gamma I-\frac{\gamma^{2}\Phi_{X}(t)\Phi_{X}(t)^{\top}}{2\|\Phi_{X}(t)\|_{F}^{2}}\geq\varepsilon I.$ (37)
@@
$v^{\top}\!\left(\gamma I-\frac{\gamma^{2}\Phi_{X}(t)\Phi_{X}(t)^{\top}}{2\|\Phi_{X}(t)\|_{F}^{2}}\right)v$ $\geq\gamma|v|^{2}-\frac{\gamma^{2}\|\Phi_{X}(t)\|_{2}^{2}}{2\|\Phi_{X}(t)\|_{F}^{2}}|v|^{2}$
$\geq\gamma|v|^{2}-\frac{\gamma^{2}}{2}|v|^{2}\geq\varepsilon|v|^{2}.$
@@
From (21), there exists a c>0c\!>\!0 such that $\operatorname{tr}\!\left(\Phi_{X}(t)^{\top}\Phi_{X}(t)\right)\!\leq\!c$ for all $t>T^{*}$. Therefore,
@@
$\lim\limits_{t\to\infty}\begin{bmatrix}A_{\rm s}+B_{\rm s}\hat{K}(t)&B_{\rm s}\hat{L}(t)\end{bmatrix}$ (42)
@@
Existing adaptive control methods with relaxed requirements imposed on data like [lavretsky2009combined, chowdhary2010concurrent, chowdhary2013concurrent, cho2017composite, lee2019concurrent, roy2017combined, katiyar2022initial] identify conditions weaker than PE: yet, only sufficiency of such conditions is proven. The key challenge solved by Theorem 5.12 is to provide a necessary and sufficient condition. The sufficiency part of the proof of Theorem 5.12 shows that
$\operatorname{im}\begin{bmatrix}I&0\\ A_{\rm m}&B_{\rm m}\end{bmatrix}\subseteq\operatorname{im}\begin{bmatrix}X_{-}(T^{*})\\ X_{+}(T^{*})\end{bmatrix}$, (43)
implies (34) and thus the asymptotic convergence in (10). Then, the proof of Theorem 5.12 also shows the necessity of (43), where Lemma 5.7 serves to connect the notion of informative data with the matching equations.
Hence, Theorem 5.12 establishes—for the first time—a necessary and sufficient condition for the asymptotic convergence of the adaptive gains to a solution of the matching equations. Notably, the proposed MRAC approach requires neither unique system identification nor structural assumptions on the system apart from controllability. We refer the reader to Section V-D for a discussion on structural assumptions in existing MRAC approaches, such as square assumptions on $L$ ($p=m$), or $B_{{\rm s}}$ being known, or $B_{{\rm s}}$ being full column rank. Before this, we discuss the boundedness of the closed-loop signals and the convergence of the error.
@@ 5.3 Stability and boundedness of the closed-loop signals
([zhou2017asymptotic]). Consider a linear time-varying autonomous system
@@
([jiang2001input]). Consider a nonlinear system
@@
Since $\lim_{t\to\infty}A_{\rm s}+B_{\rm s}\hat{K}(t)=A_{\rm m}$ and $A_{\rm m}$ is Schur, [zhou2017asymptotic, Corollary 3] implies that (47) is uniformly exponentially stable. From Definition 5.14, we have that for any $t_{0}\in\mathbb{Z}_{+}$ and $t>t_{0}$, there exist scalars $\rho^{\prime}>0$ and $\lambda^{\prime}\in(0,1)$ such that
@@
Finally, we demonstrate the convergence of the error. According to [jiang2001input, Sect. 3.2], for the input-to-state stable system (49), there exists a $\mathcal{K}$-function $f^{\prime}$ such that for any $e(0)\in\mathbb{R}^{n}$,
@@ 5.4 Discussion on excitation and structural assumptions in state-of-the-art MRAC
We now discuss the requirements on collected data imposed by the proposed MRAC method as compared to existing state-of-the-art MRAC methods. Without going into full details, we mention that state-of-the-art approaches like concurrent learning or composite MRAC [lavretsky2009combined, chowdhary2010concurrent, chowdhary2013concurrent, cho2017composite, lee2019concurrent, roy2017combined] are based on a Lyapunov design in continuous time, based on the error dynamics
@@
with $A_{\rm m}$ the Hurwitz state matrix of the reference model, $B_{\rm s}$ the control matrix of the system, $\psi$ measured signals, and $\tilde{\Lambda}$ a parameter estimation error. In this paper, we have proposed a discrete-time MRAC method. Although a direct comparison to the state-of-the-art is thus not possible, some considerations can still be made based on the analogy between the error dynamics in (49) and (51). To this end, note that all state-of-the-art solutions with the exception of [roy2017combined] require the input matrix $B_{\rm s}$ to be known, which is not required in the MRAC method in the current paper. The estimation of $B_{\rm s}$ in [roy2017combined] leveraged the concept of initial excitation to relax the classical persistence of excitation conditions. When written in discrete time, initial excitation requires the existence of $T\in\mathbb{N}$ and $\delta>0$ such that
@@
satisfy the full rank condition (4). Thus, from Lemma 2, we conclude that the requirements imposed on collected data by initial excitation are such that $(A_{\rm s},B_{\rm s})$ can be uniquely identified. Such an observation further illustrates the significance of Proposition 4.2 and Theorem 4.4. Similar to [roy2017combined], let us assume that the matching equations (6) have a solution and $p=m$. From Proposition 4.2, we know that (52) implies data informativity for model reference control. However, when $B_{\rm m}$ does not have full column rank, Theorem 4.4 implies that the initial excitation (52) is sufficient but not necessary to solve the model reference control problem. This means that initial excitation imposes stronger conditions on collected data, not required in the proposed MRAC approach.
Let us finally note that most approaches [tao2014multivariable, lavretsky2009combined, chowdhary2010concurrent, chowdhary2013concurrent, cho2017composite, roy2017combined, lee2019concurrent, katiyar2022initial] in the multi-variable MRAC literature require $B_{\rm s}$ to have full column rank, which in turns guarantees a unique solution $(K,L)$ of the matching equations (6). This is not required in the proposed MRAC method, which remains valid even when an infinite number of solutions of the matching equations (6) exist. This will be illustrated in the numerical example in the next section.